1- Stock A has a current price of $25.00, a beta of 1.25, and a dividend yield of 6%. If the Treasury bill yield is 5% and the market portfolio is expected to return 16%, what should Stock A sell for at the end of an investor’s two year investment horizon? (Hint: Solve for the growth rate using the Gordon Growth Model).

Question options:

$31.00

$31.78

$32.15

,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,

2-HMTV has planned on diversifying into the dual-PVR field. As a result, HMTV’s beta would rise to 1.6 from 1.2 and the expected future long-term growth rate in the firm’s earnings would increase from 12% to 16%. The expected market return is 14%; the risk-free rate is 7%, and the current dividend is $0.50. Should HMTV undertake the expansion?

Question options:

Yes, the stock price is expected to increase $9.89

No, the stock price is expected to decrease $3.38

No, the stock price is expected to decrease $8.28
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

ke Microsoft Excel Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy ,_a. ars-困Merge & Center, $, % , 弼,8 C Paste B 1 u. Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Font Alignment Number Cells Edting K242 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 4HKE CAPM UTILITY, SH 1 ke = Rf + beta( Rm-Rf) ke-5% + 1.25(16%-5%) ke-18.75% ke D1/P0+g 18.75% = 6% + g 8-12.75% (D1/P0- DIVIDEND YIELD-6% GIVEN) ANS PRICE OF SHARE AFTER 2 YEARS 25*(1+8)A2 25*(1+0.1275) 2 31.78 2 ke Rf + beta(Rm-Rf) ke-7% + 1.2(14%-7%)-15.4% NOW g-1296, DO-0.50 D1 DO (1+g)-0.50 (1+0.12)-0.56 P0-D1/(ke-g) = 0.56/(0.154-0.12)- 16.47 M port future INDEX INTL CAP BUD LEASING PV, FV, ANNUITY DIR beta bond c cleanYIELD bond stru WACC ex div DN11 erences: E232 130% 07:32 10-10-2018